# Pragmatic Reasoning improves LLM Code Generation

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive potential in translating natural language (NL) instructions into program code. However, user instructions often contain inherent ambiguities, making it challenging for LLMs to generate code that accurately reflects the user’s true intent. To address this challenge, researchers have proposed approaches that produce multiple candidates of the program code and then rerank them to identify the best solution. In this paper, we propose CodeRSA, a novel code candidate reranking mechanism built upon the Rational Speech Act (RSA) framework, designed to guide LLMs toward more comprehensive pragmatic reasoning about user intent. We evaluate CodeRSA using Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct on two widely used code generation benchmarks, HumanEval and MBPP. Our experiment results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance. These findings underscore the effectiveness of integrating pragmatic reasoning into code candidate reranking, offering a promising direction for enhancing code generation quality in LLMs.
@@ 1 Introduction
Recent advances in generative large language models (LLMs) have demonstrated their impressive ability to generate program code from user-provided natural language instructions (Liu et al., 2024b; Coignion et al., 2024). However, given the intrinsic complexities of coding and the potential ambiguities in user input, producing code in a single attempt may fail to explore the vast solution space, overlooking correct or higher-quality solutions (Liu et al., 2024a). A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as code candidates (Chen et al., 2021; Brown et al., 2024), and to rerank them. Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches. Due to the unreliability of automatically generated test suites (Chen et al., 2022) and the potential safety risks associated with code execution (Yetiştiren et al., 2023), we focus instead on content-driven methods, which evaluate the generated text, often relying on token-level probabilities. For example, Coder Reranking scores each candidate based on the cumulative probability of its tokens, sometimes however favoring ‘degenerate solutions” (generic or repetitive code) with disproportionately high token probabilities (Zhang et al., 2023a).
@@
When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), Coder Reranking evaluates candidate solutions solely from the listener’s perspective. Yet, research on human communication suggests that effective listeners reason about the speaker (who in turn reasons about the listener) Grice (1975). Frank and Goodman (2012; 2016) provided a principled method for quantifying this process based on a probabilistic framework rooted in game-theoretic notions, called the Rational Speech Act (RSA) framework. Pu et al. (2020, 2024) demonstrated the effectiveness of the RSA framework for program generation in a simple domain (regular-expression synthesis), while Schuster et al. (2024) reported negative results on a spreadsheet domain. One aspect that has held back RSA models from scaling up to realistic use cases is the computational overhead (Pu et al., 2024): It requires reasoning about a whole set of alternative instructions that the speaker could have given and about the set of alternative pieces of code that could solve the problem, which is computationally expensive. Zhang et al. (2023a) therefore proposed CoderReviewer Reranking as a scalable approach that simplifies these probability estimation processes over alternatives. However, this comes at the cost of not fully modelling the dialogic, interactive reasoning that can emerge when speaker and listener exchange information.
@@
We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori et al., 2024), and Qwen-2.5-7B-Instruct, a recent instruction-tuned model from the Qwen series (Yang et al., 2024), on two widely used code generation benchmarks: OpenAI’s HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). Our experimental results demonstrate that CodeRSA reliably outperforms the Coder and CoderReviewer Reranking methods. Our qualitative analysis reveals how the CodeRSA enables better candidate selection, promoting a more comprehensive understanding of user intent.
@@ 2 Related Work
Natural Language to Code. Previous research has extensively explored generating code from natural language using neural network models (Ling et al., 2016; Rabinovich et al., 2017; Hayati et al., 2018). Recently, large language models (LLMs) have propelled significant advances in this area, driven by the transformer Vaswani (2017) architecture and large-scale pretraining. Their performance on code generation tasks often surpasses that of traditional models, and in many cases even rivals human programmers (Ni et al., 2024; Becker et al., 2023). A recent study shows that LLMs also exhibit strong performance in code summarization, effectively translating code snippets into text (Akib et al., 2024).
Code Reranking Methods. Execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) evaluate code candidates by running them against automatically generated test suites. Although often effective, these methods rely on the availability and reliability of test suites, which are frequently incomplete or difficult to construct, and executing untrusted code can pose safety risks (Yetiştiren et al., 2023; Khoury et al., 2023). In contrast, content-driven reranking methods are far more versatile because they do not rely on execution and are not even confined to coding tasks.
Coder Reranking. Chen et al. (2021) rerank code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction. This process can also be called Coder Reranking because the LLM is a mere Coder that estimates the candidate probability based on the corresponding instruction. When using an LLM to estimate conditional probabilities, we compute the probability of each token iteratively. For example, in Coder Reranking, the model processes a candidate’s tokens from left to right: At each step, it calculates the probability of the current token given the instruction and the previously generated tokens, then appends that token to the context before moving on. The product of these sequential probabilities across all tokens yields the overall probability of the code candidate under the given instruction:
@@ 3 CodeRSA
where $\text{P}_{\text{LLM}}(c\mid i)$ is the probability assigned by the LLM to candidate $c$ given instruction $i$. In an idealized RSA setting, the literal listener would evaluate all possible programs, but since the space of programs is unbounded, we approximate it by sampling a finite set of candidate codes from the LLM.
We additionally define a candidate prior distribution obtained by querying the LLM without any instruction context:
@@
Clustering Paraphrases. While the basic RSA formulation operates directly over the instruction set $I$, it can suffer from over-interpreting superficial variations in wording when many instructions are semantically equivalent and differ only in surface form. In such cases, RSA allocates probability mass across paraphrases as if they were meaningful distinct alternatives, diluting the signal and reducing accuracy.
@@
To mitigate this, CodeRSA employs a semantic clustering. Candidate instructions are grouped into semantic clusters $\mathcal{C}=\{C_{1},\ldots,C_{K}\}$ using an LLM-based equivalence test (implementation details in Section 4), so that pragmatic reasoning operates over clusters rather than individual instructions. This ensures that comparisons emphasize genuine differences in meaning rather than superficial variation.
@@ 4.1 Dataset and Base Models
We evaluate on two widely used code generation benchmarks. HumanEval (Chen et al., 2021) contains 164 Python programming problems, each presented as an unfinished function with a natural language instruction. MBPP (Austin et al., 2021) includes 257 short programming tasks with natural language prompts. HumanEval offers balanced difficulty, while MBPP introduces greater lexical variety. Note that simpler datasets such as CoNaLa (Yin et al., 2018) already yield near-perfect results, leaving little room for reranking, whereas more challenging datasets such as BigCodeBench (Zhuo et al., 2024) contain many instances that cannot yet be solved by today’s state-of-the-art models, which makes it difficult to obtain meaningful comparisons of reranking methods and may obscure the performance differences we aim to study.
@@
For our experiments, we use Llama-3-8B-Instruct (Grattafiori et al., 2024) and Qwen-2.5-7B-Instruct (Yang et al., 2024), two instruction-tuned LLMs of comparable scale. Llama-3-8B-Instruct balances efficiency with strong generation quality, while Qwen-2.5-7B-Instruct provides a competitive open-source alternative. Both achieve competitive performance on HumanEval and MBPP, making them suitable for assessing reranking in our setting. We do not include specialized coder models in this study, since our framework requires both code generation and instruction-level reasoning. Future work could explore hybrid setups, for example, using coder models for program synthesis combined with general-purpose instruction-tuned models for reasoning about instructions.
@@ 5.1 Quantitative Analysis
We report further results covering different models (Qwen2.5-7B-Instruct) and datasets (HumanEval) in Appendix A.2, which confirm the same overall trends.
@@ 6 Discussion
In Section 4, we argued that CodeRSA, as a reranking approach, is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, such that we can still obtain a high quality probability distribution over instructions and over code candidates. This raises the question of the relevance of pragmatic reasoning for code generation, and more generally in communication. Research on human communication has demonstrated the importance of pragmatic reasoning, even though it introduces additional computational overhead. At the same time, studies suggest that humans may rely on simple heuristics or amortized estimates (Pu et al., 2024), avoiding iterative reasoning in easy cases while still engaging in full pragmatic reasoning when tasks are more complex.
@@ 7 Conclusion
This work introduces CodeRSA, a candidate reranking algorithm for the generation of program code grounded in the Rational Speech Act framework. By modeling the iterative reasoning of a pragmatic listener about a pragmatic speaker, CodeRSA consistently outperforms the Coder Reranking baseline and surpasses the state-of-the-art CoderReviewer approach. A qualitative analysis further reveals that, even when incorporating certain idealized assumptions and variations, CodeRSA remains faithful to the core principles of the RSA framework. These results highlight the effectiveness of applying well-established linguistic frameworks to enhance reasoning in language models, opening new avenues for research and development in code-related tasks.
@@ 8 Limitations
Another limitation is that, although our experiments already cover two models (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) and two datasets (HumanEval and MBPP), the scope remains relatively narrow. We are currently working on incorporating additional balanced-difficulty datasets such as DS-1000 (Lai et al., 2023)), along with further open-source models like Mistral (Jiang et al., 2023) and newer Qwen releases beyond Qwen-2.5. This expansion will allow us to evaluate reranking methods across a wider range of scenarios, ultimately leading to a more comprehensive assessment of our approach.
@@ Appendix A Appendix
A.1 A Conjecture: Explaining CoderReviewer from an RSA Perspective
In the RSA framework, a pragmatic listener’s posterior over a candidate $c$ given an instruction $i$ is commonly expressed as:
$\text{P}_{L_{1}}(c\mid i)\propto\text{P}_{S_{1}}(i\mid c)\cdot\text{P}(c),$
where $\text{P}_{S_{1}}(i\mid c)$ represents how likely a pragmatic speaker would be to produce instruction $i$ when the correct candidate is $c$, and $\text{P}(c)$ is the prior likelihood of $c$.
Translating this perspective to LLMs, we hypothesize that when generating instructions (the “Reviewer” role), it is relatively straightforward for the model to produce abstract instructions from concrete code. Since code is unambiguous, the LLM can approximate a pragmatic speaker:
$\text{P}_{\text{LLM}}(i\mid c)\approx\text{P}_{S_{1}}(i\mid c).$
However, generating code from abstract instructions (the “Coder” role) is substantially more difficult. In this setting, the LLM may effectively revert to estimating a prior over possible candidates, thereby approximating:
$\text{P}_{\text{LLM}}(c\mid i)\approx\text{P}(c).$
From this RSA standpoint, the CoderReviewer paradigm can be considered a simplified, yet broad, modeling of a pragmatic listener.
@@ A.2 More Details of Results
A.3 Accuracy vs. Number of Sampled Candidates (MBPP)

### v4

@@ Abstract
Pragmatic reasoning is pervasive in human–human communication — it allows us to leverage shared knowledge and counterfactual reasoning in order to infer the intention of a conversational partner given their ambiguous or underspecified message. In human–computer communication, underspecified messages often represent a major challenge: for instance, translating natural language instructions into code is difficult when user instructions contain inherent ambiguities. In the present paper, we aim to scale up the pragmatic “Rational Speech Act” framework to naturalistic language-to-code problems, and propose a way of dealing with multiple meaning-equivalent instruction alternatives, an issue that does not arise in previous toy-scale problems. We evaluate our method, CodeRSA, with two recent LLMs (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) on two widely used code generation benchmarks (HumanEval and MBPP). Our experimental results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance. Qualitative analyses demonstrate that it exhibits the desired behaviour for the right reasons. These findings underscore the effectiveness of integrating pragmatic reasoning into a naturalistic complex communication task, language-to-code generation, offering a promising direction for enhancing code generation quality in LLMs and emphasizing the importance of pragmatic reasoning in complex communication settings.
@@ 1 Introduction
Recent advances in generative large language models (LLMs) have demonstrated their impressive ability to generate program code from user-provided natural language instructions (Liu et al., 2024b; Coignion et al., 2024). However, given the intrinsic complexities of coding and the potential ambiguities in user input, producing code in a single attempt may fail to explore the vast solution space, overlooking correct or higher-quality solutions (Liu et al., 2024a). A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as code candidates (Chen and others, 2021; Brown et al., 2024), and to rerank them.
@@
When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), the different code candidates represent the different possible user intents and the instruction represents an ambiguous or underspecified message. Pragmatic reasoning then involves considering not only the different possible intended meanings, but also counterfactual reasoning about what instruction the speaker would have given if they had a specific intent. These ideas go back to Grice (1975) and have been mathematically formalized using a probabilistic model rooted in game-theoretic notions, as the "Rational Speech Act" (RSA) framework in Frank and Goodman (2012; 2016). Code reranking can thus be performed with respect to the RSA reasoning process. In previous research, Pu et al. (2020, 2024) demonstrated the effectiveness of the RSA framework for program generation in a simple domain (regular-expression synthesis), while Schuster et al. (2024) reported negative results on a spreadsheet domain. One aspect that has held back RSA models from scaling up to realistic use cases is the computational overhead (Pu et al., 2024): It requires reasoning about a whole set of alternative instructions that the speaker could have given and about the set of alternative pieces of code that could solve the problem, which is computationally expensive. Zhang et al. (2023a) therefore proposed CoderReviewer Reranking as a scalable approach that simplifies these probability estimation processes over alternatives. However, this comes at the cost of not fully modelling the dialogic, interactive reasoning that can emerge when speaker and listener exchange information.
@@
We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori and others, 2024), and Qwen-2.5-7B-Instruct, a recent instruction-tuned model from the Qwen series (Yang et al., 2024), on two widely used code generation benchmarks: OpenAI’s HumanEval (Chen and others, 2021) and MBPP (Austin et al., 2021). Our experimental results demonstrate that CodeRSA reliably outperforms the simpler Coder and CoderReviewer Reranking methods. Our qualitative analysis reveals how the CodeRSA enables better candidate selection, promoting a more comprehensive understanding of user intent.
@@ 2 Related Work
Natural Language to Code. Previous research has extensively explored generating code from natural language using neural network models (Ling et al., 2016; Rabinovich et al., 2017; Hayati et al., 2018). Recently, large language models (LLMs) have propelled significant advances in this area, driven by the transformer Vaswani et al. (2017) architecture and large-scale pretraining. Their performance on code generation tasks often surpasses that of traditional models, and in many cases even rivals human programmers (Ni et al., 2024; Becker et al., 2023). A recent study shows that LLMs also exhibit strong performance in code summarization, effectively translating code snippets into text (Akib et al., 2024).
Code Reranking Methods.
Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches. Execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) evaluate code candidates by running them against automatically generated test suites. Although often effective, these methods rely on the availability and reliability of test suites, which are frequently incomplete or difficult to construct, and executing untrusted code can pose safety risks (Yetiştiren et al., 2023; Khoury et al., 2023). In contrast, content-driven reranking methods are far more versatile because they do not rely on execution and are not even confined to coding tasks. One such model is Coder Reranking (Chen and others, 2021).
Coder Reranking.
Chen and others (2021) rerank code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction. This process can also be called Coder Reranking because the LLM is a mere Coder that estimates the candidate probability based on the corresponding instruction. When using an LLM to estimate conditional probabilities, we compute the probability of each token iteratively. For example, in Coder Reranking, the model processes a candidate’s tokens from left to right: At each step, it calculates the probability of the current token given the instruction and the previously generated tokens, then appends that token to the context before moving on. The product of these sequential probabilities across all tokens yields the overall probability of the code candidate under the given instruction:
@@ 3 CodeRSA
where $\text{P}_{\text{LLM}}(c\mid i)$ is the probability assigned by the LLM to candidate $c$ given instruction $i$. In an idealized RSA setting, the literal listener would evaluate all possible programs, but since the space of programs is unbounded, we approximate it by sampling a finite set of candidate codes from the LLM. We additionally define a candidate prior distribution obtained by querying the LLM without any instruction context:
@@
Clustering Paraphrases. While the basic RSA formulation operates directly over the instruction set $I$,
@@
it can suffer from over-interpreting superficial variations in wording when many instructions are semantically equivalent and differ only in surface form. In such cases, RSA allocates probability mass across paraphrases as if they were meaningful distinct alternatives, diluting the signal and reducing accuracy.
To mitigate this, CodeRSA employs semantic clustering. Candidate instructions are grouped into semantic clusters $\mathcal{C}=\{C_{1},\ldots,C_{K}\}$ using an LLM-based equivalence test (implementation details in Section 4), so that pragmatic reasoning operates over clusters rather than individual instructions. This ensures that comparisons emphasize genuine differences in meaning rather than superficial variation.
@@ 4.1 Dataset and Base Models
We evaluate on two widely used code generation benchmarks. HumanEval (Chen and others, 2021) contains 164 Python programming problems, each presented as an unfinished function with a natural language instruction. MBPP (Austin et al., 2021) includes 974 short programming tasks with natural language prompts. HumanEval offers balanced difficulty, while MBPP introduces greater lexical variety. Note that simpler datasets such as CoNaLa (Yin et al., 2018) already yield near-perfect results, leaving little room for reranking, whereas more challenging datasets such as BigCodeBench (Zhuo et al., 2024) contain many instances that cannot yet be solved by today’s state-of-the-art models, which makes it difficult to obtain meaningful comparisons of reranking methods and may obscure the performance differences we aim to study.
@@
For our experiments, we use Llama-3-8B-Instruct (Grattafiori and others, 2024) and Qwen-2.5-7B-Instruct (Yang et al., 2024), two instruction-tuned LLMs of comparable scale. Llama-3-8B-Instruct balances efficiency with strong generation quality, while Qwen-2.5-7B-Instruct provides a competitive open-source alternative. Both achieve competitive performance on HumanEval and MBPP, making them suitable for assessing reranking in our setting. We do not include specialized coder models in this study, since our framework requires both code generation and instruction-level reasoning. Future work could explore hybrid setups, for example, using coder models for program synthesis combined with general-purpose instruction-tuned models for reasoning about instructions.
Our goal is to examine whether RSA-style pragmatic reasoning can be effectively applied to realistic code generation through appropriate methodological design. We evaluate our approach on two established benchmarks (MBPP and HumanEval) using mid-sized instruction-tuned models, which provide sufficient headroom for reranking. For example, Llama-3-8B-Instruct improves from 51.3% pass@1 to 78.2% pass@10 on MBPP, indicating that correct solutions are often generated but not reliably selected in a single attempt. In such settings, the effect of reranking is most clearly observable. Appendix A.3 further reports a sanity check with Llama-3-70B-Instruct Grattafiori and others (2024), confirming that CodeRSA behaves consistently when applied to a substantially larger model.
@@ 5.1 Quantitative Analysis
We report further results covering different models (Qwen2.5-7B-Instruct) and datasets (HumanEval) in Appendix A.2, which confirm the same pattern of CodeRSA consistently improving performance over baseline methods.
@@ 6 Discussion
In Section 4, we argued that CodeRSA is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, i.e., good candidates need to be contained in the set of sampled candidates which is then re-ranked. That is, an important pre-condition for the applicability of the CodeRSA approach is the ability to obtain a high quality probability distribution over instructions and over code candidates. We argue that in complex tasks like language instructions for generating code, a situation where the correct candidate is in the set of sampled candidates, but where it is not ranked first, is not just due to bad model performance that will become better with improved models, but is expected to be a systematic phenomenon that is due to the inherent ambiguity and underspecification in human natural language instructions. Dealing with this type of ambiguity is not a matter of better language modelling, but inherently requires the counterfactual reasoning process implemented in RSA models. Research on human communication has demonstrated the importance of pragmatic reasoning, even though it introduces additional computational overhead. At the same time, studies suggest that humans may switch between different strategies where they sometimes rely on the expensive reasoning process, and sometimes rely on simple heuristics or amortized estimates (Lieder and Griffiths, 2020; Tscshantz et al., 2023; Dasgupta et al., 2018; Pu et al., 2024), avoiding iterative reasoning in easy cases while still engaging in full pragmatic reasoning when tasks are more complex. Such strategies can also be explored in future CodeRSA research.
@@ 7 Conclusion
This work introduces CodeRSA, a pragmatics-inspired approach to dealing with ambiguity and underspecification in human instructions for code generation. By modeling the iterative reasoning of a pragmatic listener about a pragmatic speaker, CodeRSA consistently outperforms the Coder Reranking baseline and surpasses the state-of-the-art CoderReviewer approach. Our experiments showed highly consistent results for two models from different model families and two different benchmarks. A qualitative analysis further reveals that, even when incorporating certain idealized assumptions and variations, CodeRSA remains faithful to the core principles and goals of the RSA framework. These results highlight the effectiveness of applying well-established linguistic frameworks to enhance reasoning in language models, opening new avenues for research and development in code-related tasks. Some of our insights related to practical and naturalistic application domains, such as the need for clustering of very similar instructions or the need to re-weight probability estimates for instructions vs. meaning representations may also be applicable to human-computer interaction tasks beyond language-to-code.
@@ 8 Limitations
Another limitation concerns the scope of models and benchmarks considered. In the current study, we deliberately focus on mid-sized models and well-established benchmarks as a proof-of-concept design choice, enabling fair and clearly observable comparisons with existing reranking baselines, rather than aiming for exhaustive coverage across model scales or benchmark families. As a result, although our experiments already cover two models (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) and two datasets (HumanEval and MBPP), the overall scope remains relatively narrow. In future work, we plan to incorporate additional balanced-difficulty datasets such as DS-1000 (Lai et al., 2023), along with further open-source models like Mistral (Jiang et al., 2023) and newer Qwen releases beyond Qwen-2.5. This expansion will allow us to evaluate reranking methods across a wider range of scenarios, ultimately leading to a more comprehensive assessment of our approach. At the same time, ensuring an appropriate match between model capability and benchmark difficulty remains an important challenge for designing informative evaluations.
@@ Appendix A Appendix
A.1 A Conjecture: Viewing CoderReviewer through an RSA Lens
CoderReviewer reranks candidates using the bidirectional criterion
$\text{P}_{\text{CR}}(c\mid i_{0})\ \propto\ \text{P}_{\text{LLM}}(c\mid i_{0})\cdot\text{P}_{\text{LLM}}(i_{0}\mid c),$
which rewards candidates that are (i) likely under the instruction (Coder) and (ii) able to “explain” the instruction when the code is given as context (Reviewer).
In RSA, a pragmatic listener instead ranks candidates according to
$\text{P}_{L_{1}}(c\mid i_{0})\ \propto\ \text{P}_{S_{1}}(i_{0}\mid c)\cdot\text{P}(c),$
combining (a) a speaker term that captures how well the instruction identifies the intended candidate among alternatives, and (b) a prior over candidates.
We conjecture that CoderReviewer can be interpreted as a lightweight, heuristic approximation to this RSA listener objective on the restricted candidate set used for reranking. Intuitively, producing a faithful natural-language description from a concrete piece of code is comparatively constrained, so the reverse model $\text{P}_{\text{LLM}}(i_{0}\mid c)$ may behave like an amortized proxy for a speaker preference $\text{P}_{S_{1}}(i_{0}\mid c)$ (without explicitly normalizing over an alternative-instruction set or modeling instruction costs). Meanwhile, when $i_{0}$ is underspecified, the forward term $\text{P}_{\text{LLM}}(c\mid i_{0})$ can be strongly influenced by instruction-independent code preferences (well-formedness, idiomatic patterns, and length/format biases), and thus acts as a task-local proxy for the candidate prior $\text{P}(c)$ on the finite set of sampled candidates. Under this view, multiplying the two terms resembles combining a speaker-like compatibility signal with a prior-like plausibility signal, as in RSA.
@@ A.2 More Details of Results
A.3 Additional Experiment: Larger-Model Sanity Check on MBPP
As an additional sanity check, we evaluate CodeRSA with a larger base model. Since RSA-style reranking is computationally expensive, we keep this experiment limited in scope. We use Llama-3-70B-Instruct Grattafiori and others (2024) on MBPP and restrict evaluation to instances that are unsolved at pass@1 but solved at pass@10, resulting in 84 problems where reranking is non-trivial.
Method Solved Acc.
Coder 33/84 39.3%
CoderReviewer 41/84 48.8%
CodeRSA (best $\alpha=0.9$) 43/84 51.2%
CodeRSA (avg. $\alpha\in[0.5,1.5]$) 35.5/84 42.3%
Table 2: Sanity check on Llama-3-70B-Instruct for MBPP, evaluated on 84 instances that are unsolved at pass@1 but solved at pass@10. For CodeRSA, “best” is the best value in a sweep with step size 0.1; “avg.” is the mean over $\alpha\in\{0.5,0.6,\ldots,1.5\}$.
Table 2 reports the number of solved problems (out of 84) and the corresponding accuracy after reranking. CodeRSA achieves the best result at $\alpha=0.9$ in our sweep (step size 0.1), outperforming Coder and slightly improving over CoderReviewer on this subset. Averaged over $\alpha\in[0.5,1.5]$, CodeRSA remains above Coder, indicating that scaling to a larger model does not lead to abnormal behavior, while calibration still matters for peak performance.
A.4 Accuracy vs. Number of Sampled Candidates (MBPP)